sally can paint a room in 3 hours white it takes steve 4 hours to paint the same room. How long would it take them to paint the room if they worked together?
Answer by richwmiller(17219)

x/3+x/4=1
or
1/3+1/4=1/x
4/12+3/12=1/x
7/12=1/x
12/7=x